OpenStudy (lovelyharmonics):

ellipses c:

OpenStudy (lovelyharmonics):

Find the center, vertices, and foci of the ellipse with equation x squared divided by one hundred plus y squared divided by sixty four = 1

ganeshie8 (ganeshie8):

use the attached chart

ganeshie8 (ganeshie8):

\(\dfrac{x^2}{100} + \dfrac{y^2}{64} = 1\) is same as : \(\dfrac{x^2}{10^2} + \dfrac{y^2}{8^2} = 1\)

ganeshie8 (ganeshie8):

comparing it wid standard form : \(\dfrac{x^2}{a^2} + \dfrac{y^2}{b^2} = 1\) gives : \(a = 10, ~b = 8\)

ganeshie8 (ganeshie8):

see if u can use the chart to find out vertices/center/foci

OpenStudy (lovelyharmonics):

how do you find the center? @ganeshie8

OpenStudy (lovelyharmonics):

and also, how do you find c? @ganeshie8

ganeshie8 (ganeshie8):

center is (0, 0) for these ellipses...

ganeshie8 (ganeshie8):

these ellipses are centered at origin

ganeshie8 (ganeshie8):

use this for finding "c" : c^2 = a^2 - b^2

OpenStudy (lovelyharmonics):

so 100- 64= 36 which would be 6 okay so the verticies are -10,0 and 10,0 and the foci are -6,0 and 6,0 c: @ganeshie8

ganeshie8 (ganeshie8):

Excellent !

OpenStudy (lovelyharmonics):

yay ^.^ if i post more questions like this will you help on thoses too c:

ganeshie8 (ganeshie8):

sure :)

ganeshie8 (ganeshie8):

but the attached chart is all u need... I see you're getting the hang of it :)

OpenStudy (lovelyharmonics):

yay c: hold on how do you tell the difference of whether they are vertical or eliptical because the only difference is the denominator, which dosent tell me whether it is a or b

ganeshie8 (ganeshie8):

very good question !

ganeshie8 (ganeshie8):

whichever is largest is the \(\large a\)

ganeshie8 (ganeshie8):

\(\large a\) = length of semi-major axis

ganeshie8 (ganeshie8):

semi-major is axis is always alrger than the semi-minor axis

ganeshie8 (ganeshie8):

So, in short : whatever denominator is LARGE is the \(\large a\)

OpenStudy (lovelyharmonics):

ohh that makes sense c:

OpenStudy (lovelyharmonics):

and how do you know when an ellipses isnt centered at the origin?

ganeshie8 (ganeshie8):

if the center is (h, k) then the equation of ellipse wud look like : http://www.teacherschoice.com.au/images/ellipse_types.gif

OpenStudy (lovelyharmonics):

wait so this equation x squared divided by nine plus y squared divided by sixteen = 1 would be centered right?

OpenStudy (lovelyharmonics):

@ganeshie8

ganeshie8 (ganeshie8):

do u mean it is centered at origin ? then yes

ganeshie8 (ganeshie8):

any ellipse of form : \(\dfrac{x^2}{a^2} + \dfrac{y^2}{b^2} =1\) is centered at \((0,0)\)

OpenStudy (lovelyharmonics):

yay c: thank you

ganeshie8 (ganeshie8):

:)